Minutes of the Linac-Booster-ISOLDE Supervisor Meeting held on
Present: G.Bellodi, A.Findlay, F.Gerigk, K.Hanke, D.McFarlane, B.Mikulec, G.Rumolo, J.Tan
Media event for first 3.5 TeV collisions next Tuesday; expect some perturbations, and the BBC
will occupy “our” room. No risk for the machine until then!
F.Gerigk → G.Bellodi
During the technical stop last Monday/Tuesday a number of tests on the source were
done, to understand the regular intensity dips; nothing abnormal was found (looked at
magnets, cathode, gas supply, hv column was cleaned, ...).
Before beam was put back on Tuesday (16 March), access was done to check LT.QND10
and 12 which make noise; for the moment no reason to worry, magnets stay under close
Thursday (18 March) problem with LA2.QDN44, it was pulsing with wrong value and
caused a radiation alarm on PAX22, the EPC piquet changed a regulation card.
Saturday (20 March) rf fault tank1, the Booster operator reset the Franck-James.
3. PS Booster
A.Findlay → K.Hanke
Tuesday 16 March
- A blind access on door 703 caused an alarm, hence a patrol was required.
- Technical stop was due to finish at 18H00, the MPS was up and running, the cavities
all operational by 18H15 and the extraction kicker for R4 back in action by 18H30.
By 18H50 the LHCINDIV beam was being used to de-bug the machine. A few
multipoles needed a local reset by the operator, kickers needed some resets, but by
20H15 the INDIV beam was available for the PS. The BTP.QNO60 required an
intervention by the EPC piquet to try to get it back up and running. By 22H15 the
PROBE beam was ready for the PS. The BTP.QNO60 was reluctant to behave as we
expect, requiring a local intervention by the operator and advice by the Pi.CO, after
which it was left in local with the correct value programmed, until the following
Wednesday 17 March
- The piquets CO & EPC were contacted to continue to de-bug the BTP.QNO60, and
after a second change of the cpu card, it was behaving as required.
- A card was changed on DPSBOBSV by M.Catin.
- The Stray Field Compensation was enabled once more, and was working correctly.
- The EPC piquet was called to fix the BRx.QSKHO, as they had all tripped, but after
20 minutes they were repaired.
- A water leak on the tuning of the BR4.C04 was fixed by M.Haase.
- BTP.MBL00 was found to be working once more after a check by D.Allen.
Thursday 18 March
- Setting up of SFTPRO for nominal intensity was started.
- LHC75 was finished for nominal intensity.
- The LA2.QDN44 started to pulse with the wrong value causing a radiation alarm on
PAX22. The EPC piquet was called into to fix it, and 30 minutes later it was
operational once more.
Friday 19 March
- The AD was installed for setting up in the PS, as this user had already been set up in
- LHC50 was used for measurements in the PSB.
- At 23H00 the MPS started to trip, and when resets would not work, the EPC piquet
was called in. He got it back on its feet by midnight, but at 05H30 it started to trip
again, requiring another intervention by the EPC piquet. By 07H15 he thought he had
identified the problem, hence switched to the TRIM A supply and we got the beam
Saturday 20 March
- The LINAC2 RF tank 1 dropped out at midday due to the Frank James power supply,
and a local reset by the operator was required, but 35 minutes later it was back in
- By 14H30 the MPS started to play up again, and remote resets would not solve the
problem, so the EPC piquet was called back in again. The piquet discussed with the
equipment specialist and they decided to swap back to the TRIM A supply. The
operator joined the piquet in the PSB to keep up to date with the progress, and by
15H30 it was thought a solution had been found so the piquet could leave. This was
short lived however, and 15 minutes after he had left, the operator called him back as
the MPS had dropped out again. A swap back to the TRIM B supply got the MPS
back online by 16h00, but the piquet joined the operator in the CCC to keep an eye on
the supply. After an hour the MPS seemed to be happy once more, so the piquet was
allowed to leave!
- BTY.DVT212 was found not to have any signal on OASIS, so a CO issue was sent.
Sunday 21 March
- The LINAC2 EAST B setting up was started by Herve.
- At 22H30 a bad contact in an emergency stop button in the PS stopped the beam. In
the PSB this was seen due to LT.VVS20 and BTP.VVS20 closing, but by 02H00 the
PSB was back online.
Monday 22 March
- The operation display wouldn't work, so the operator called the CO piquet to reboot
the proxy server.
- LHCPROBE was used by R.Scrivens for measurements in the LINAC.
- LHC50 was used for measurements by A.Findlay.
- Work on setting up SFTPRO was continued by the operator.
Some details on the MPS problems:
- Max Chamiot-Clerc explained to me that he has a different problem on each of the
TRIM supplies, but he has identified the cause in each case and is working on
- On one of the supplies (TRIM A) there is a connector that develops a bad contact
giving a false interlock. This connector he is in the process of changing, but as a
longer term solution, he is looking into other ways of cabling the interlocks.
- The other supply (TRIM B) has a connector on the backplane of a crate that develops
a bad contact from time to time. This connector he will change, but as he expects it to
take a day, he'd prefer to do it in the shadow of a stop.
- This wasn't to be however, as at 14H00 the MPS went down again, so TRIM B was
replaced by TRIM A, and Max had to do the repair of the back plane connector on
TRIM B. We ran with TRIM A during the afternoon (with a couple of failures of this
PSU), but by 18H00 Max had repaired the TRIM B, reinstalled it and we've been
running smoothly ever since.
Beams which are fully operational in the PSB:
LHCPROBE, LHCPILOT, LHCINDIV, LHC25A&B (nominal intensity), LHC50, LHC75
(nominal intensity), AD, NORMGPS&HRS & MD1.
Beams being set up:
CNGS (final setup being done by Fabrice), SFTPRO (R3 good, done by Mohammad+OPS+sup,
others to do), EAST B (work started by Herve).
Beams still to do:
LHC25 intermediate intensities, LHC75 intermediate intensities, EAST A & EAST C, TOF
6. For the coming week
Continue setting up according to list above.
Simone requests MTE beam with 500E10 total (4 rings). Need to identify a user. If possible get
SFTPRO ready, archive it and then work on the SFTPRO user.